# Tracking Meets LoRA: Faster Training, Larger Model, Stronger Performance

### v1

@@ Abstract
Motivated by the Parameter-Efficient Fine-Tuning (PEFT) in large language models, we propose LoRAT, a method that unveils the power of larger Vision Transformers (ViT) for tracking within laboratory-level resources. The essence of our work lies in adapting LoRA, a technique that fine-tunes a small subset of model parameters without adding inference latency, to the domain of visual tracking. However, unique challenges and potential domain gaps make this transfer not as easy as the first intuition. Firstly, a transformer-based tracker constructs unshared position embedding for template and search image. This poses a challenge for the transfer of LoRA, usually requiring consistency in the design when applied to the pre-trained backbone, to downstream tasks. Secondly, the inductive bias inherent in convolutional heads diminishes the effectiveness of parameter-efficient fine-tuning in tracking models. To overcome these limitations, we first decouple the position embeddings in transformer-based trackers into shared spatial ones and independent type ones. The shared embeddings, which describe the absolute coordinates of multi-resolution images (namely, the template and search images), are inherited from the pre-trained backbones. In contrast, the independent embeddings indicate the sources of each token and are learned from scratch. Furthermore, we design an anchor-free head solely based on a multilayer perceptron (MLP) to adapt PETR, enabling better performance with less computational overhead. With our design, 1) it becomes practical to train trackers with the ViT-g backbone on GPUs with only memory of 25.8GB (batch size of 16); 2) we reduce the training time of the L-224 variant from 35.0 to 10.8 GPU hours; 3) we improve the LaSOT SUC score from 0.703 to 0.743 with the L-224 variant; 4) we fast the inference speed of the L-224 variant from 52 to 119 FPS. Code and models will be released.
@@ 1 Introduction
Yet, it is not as easy as our first intuition to apply LoRA to the one-stream tracking architecture, mainly due to two key issues below: ① Ineffective design of positional encoding module. Most Transformer-based trackers adopt separate positional encodings for template and search region tokens. Albeit effective in full fine-tuning [62], such design is not well aligned with LoRA-style PEFT and thus results in suboptimal performance. We argue that this ineffectiveness comes from the disruption to the original structure of the pre-trained model, which is critical to preserve in PEFT such as LoRA [30]. ② Inductive biases. We empirically identify the convolutional head in the one-stream tracker [62] as a bottleneck in achieving convergence with LoRA. We argue that this is caused by strong inductive biases in convolutional network [15], making it hard to adapt the pre-trained model by fine-tuning a small subset of parameters using LoRA.
Our solution. To overcome the above issues in employing LoRA to one-stream tracker [62], we present two embarrassingly simple but effective designs. Specifically, in addressing issue ①, we draw inspiration from the well-known BERT [13] and introduce the token type embedding in input embedding. Concretely, we assign specific types to tokens of template and search region. By doing so, we can directly leverage existing positional encoding modules of pre-trained ViT models without disrupting their original structure. Considering different resolutions of template and search regions, we further devise a multi-resolution absolute positional embedding adaptation strategy, enabling the support of multi-resolution inputs to pre-trained ViT models for better performance. To deal with issue ②, we propose a multilayer perceptron (MLP)-based anchor-free head network, instead of convolutional head network as in [62], for target classification and bounding box regression. Such a design effectively avoids the inductive biases when applying LoRA to our baseline one-stream tracker, enabling better performance with less computational overhead as shown in our experiments.
[…]
@@ 2.1 Paradigms of Transformer Tracker
More recently, the one-stream tracking framework emerges as a significant evolution, with Mixformer [8], SimTrack [5], and OSTrack [62] serving as the vanguard. This paradigm allows for earlier interaction between the template and search region, facilitating joint feature extraction and fusion. The self-attention mechanism of the Transformer architecture plays a crucial role in this context, enhancing the model’s ability to capture complex dependencies. Most subsequent Transformer trackers follow this paradigm. GRM [21] introduces adaptive token division to enable more flexible relation modeling. DropTrack [59] adaptively performs spatial-attention dropout in the frame reconstruction to facilitate temporal correspondence learning in videos. SeqTrack [6] and ARTrack [58] add a Transformer decoder on the top a one-stream encoder to predict object bounding boxes in an autoregressive fashion.
@@ 3 Adapting ViT for LoRA-friendly Tracker
Figure 2: Architecture of LoRAT. The template and search region are first split and then projected as patch embeddings. Patch embeddings are added with shared position embeddings and token type embeddings as the input embeddings, which are then fed into Transformer encoder for joint feature extraction and fusion. The resulting representations are fed to the MLP-only head network for target classification and anchor-free based bounding box regression. Most network components from the pre-trained ViT model are frozen during training, except for LoRA modules applied on the linear layers in the Transformer encoder, the token type embeddings, and the head network.
This section meticulously outlines our approach, beginning with an elucidation of our baseline, the one-stream tracking framework. We then introduce the concept of Low-Rank Adaptation (LoRA), a pivotal technique for the efficient tuning of large-scale models for specialized tasks, emphasizing its role in enhancing model adaptability with minimal computational overhead. Following this, we explore the architectural adjustments necessary to render our model compatible with LoRA, including modifications to the model’s input embedding and an MLP-only head network. An overview of our tracker is shown in Fig. 2.
@@ 3.1 Preliminaries
One-stream trackers follow the Siamese tracker framework [1], requiring a template image $T\in\mathbb{R}^{H_{T}\times W_{T}\times 3}$ and a search region image $S\in\mathbb{R}^{H_{S}\times W_{S}\times 3}$ as inputs. The template image $T$ and the search region image $S$ are first projected by a learnable linear projection layer to obtain the template token embeddings and the search region token embeddings. Assuming the image patch resolution is $(P,P)$, the template image is divided into $N_{T}=N_{T}^{H}\times N_{T}^{W}$ image patches, and the search region image is divided into $N_{S}=N_{S}^{H}\times N_{S}^{W}$ image patches, where ${N_{T}^{H},N_{T}^{W},N_{S}^{H},N_{S}^{W}}={H_{T},W_{T},H_{S},W_{S}}/P$. The resulting patch embeddings are the template token embeddings $\mathbf{T}\in\mathbb{R}^{N_{T}\times d}$ and the search region embeddings $\mathbf{S}\in\mathbb{R}^{N_{S}\times d}$, where $d$ represents the hidden dimension of the network.
@@
Low-Rank Adaptation (LoRA) [30] is a technique for the efficient fine-tuning of large-scale pre-trained models. Typically, these models are initially trained on extensive datasets, which provides a broad foundational knowledge. To adapt the models to specific downstream tasks, models are fine-tuned with domain-specific data. Traditional fine-tuning is computationally intensive, as it involves a full forward and backward pass through the network, requiring significant computational resources and memory.
LoRA addresses these issues by proposing a low-rank matrix decomposition to represent weight updates, which can hugely reduce the number of trainable weights and memory requirements. The principle behind LoRA is that while neural networks typically have full-rank weight matrices during pre-training, the updates required for fine-tuning to a specific task have a low “intrinsic dimension”. Remarkably, in many applications, LoRA shows a comparable performance to full fine-tuning of the network.
@@
Importantly, adopting LoRA does not increase the inference latency of the model. After fine-tuning, the model’s weights are updated by adding the low-rank approximations to the original weights. This approach simplifies the deployment of multiple task-specific models on top of a single large model because the updates ($\Delta\Phi$) are much smaller than the original weights ($\Delta\Theta$).
@@ 3.2 LoRA-friendly Model Design
ViT models require positional encodings to inject positional information. Existing one-stream trackers use separate positional embeddings for template and search region tokens, which disrupt the original structure of the pre-trained ViT model, leading to ineffectiveness in PEFT. We explore multi-resolution positional embedding and token type embedding to address this issue.
@@
Token type embedding. We incorporate token type embedding (also called segment embedding), which is originally proposed in the BERT[13] language model, into our tracker, assigning unique embeddings for template and search region tokens. This approach decouples token type identification from positional embeddings, improving parameter efficiency and flexibility.
@@
We also explore more usage of token type embedding. The cropping strategy [1] commonly used by Siamese trackers to generate tracking templates has several issues that may confuse the tracker: the aspect ratio of the target objects in the templates is variable, some target objects do not have clear semantic meanings, and some target objects in the templates can not be not clearly distinguishable from the background. Token type embedding can help alleviate these issues by explicitly annotating foreground and background parts.
@@
Positional embedding. Most ViT models adapt the 1D absolute positional embedding [15], which only works for fixed resolutions. Meanwhile, the one-stream trackers usually utilize images of two different sizes as inputs. According to the PEFT-friendly model design principle, we have to share the positional embedding across two images. We explore two strategies to reuse existing absolute positional embeddings for different input sizes. Both strategies regard the original 1D absolute embedding in a 2D view. Suppose the original 1D absolute positional embedding is $P$, the 2D view of $P$ is $P_{2d}=[p_{1,1},...,p_{1,w};p_{2,1},...,p_{2,w};p_{h,1},...,p_{h,w}]$. For convenience, we assume that the resolution of search region image is the native resolution of ViT model, thus $P_{2d}$ can be directly applied to search region tokens, i.e. the search region positional embedding $P_{x}=P_{2d}$.
@@
The first approach regards absolute positional embedding as continuous spatial position embedding. In this way, we can direct interpolate the $P_{x}$ to the expected size of $P_{z}$. In this way, $P_{z}=\mathrm{interpolate}(P_{x})$. We call this approach as interpolation-based adaptation.
@@
The second approach regards absolute positional embedding as discrete patch index embedding. Thus, the smaller positional embedding for template tokens $P_{z}$ can be directly taken as a sub-matrix of the $P_{x}$, which is formed by taking a block of the entries of the size of template patches $(N_{T}^{H}\times N_{T}^{W})$ from the top left corner of the $P_{x}$. In this way, $P_{z}=[p_{1,1},...,p_{1,N^{W}_{T}};p_{2,1},...,p_{2,N^{W}_{T}};p_{N^{H}_{T},1},...,p_{N^{H}_{T},N^{W}_{T}}]$. We call this approach as slicing-based adaptation.
@@
For other commonly used positional encoding schemes, like sinusoidal positional encoding [15], relative positional embedding[52] or Rotary Positional Embedding[53], they do not target on specific resolution, so apply them on template tokens and search region tokens individually.
@@
To alleviate potential inductive biases brought by the convolutional head, such as locality assumptions on data structure [15], which may hinder convergence with LoRA-based fine-tuning, we employ a multi-layer perceptron (MLP) only head.
The head network consists of two branches for target classification and bounding box regression, each implemented as a three-layer perceptron. Both branches take search area tokens $\mathcal{S}$ from the Transformer encoder output as input. The head network applies a center-based anchor-free [54] style. The target classification branch predicts a classification confidence map $\mathbf{R}\in(0,1)^{N_{S}^{H}\times N_{S}^{W}}$, estimating the IoU between the predicted bounding box and the ground-truth (the estimation value is targeted to 0 during training if the object does not fall in the current anchor). The bounding box regression branch predicts a bounding box map $\mathbf{B}\in(0,1)^{N_{S}^{H}\times N_{S}^{W}\times 4}$. The anchor with the maximum value on the classification confidence map is considered the responsible anchor. For each anchor center point $(x_{i},y_{j})$, the bounding box regression branch predicts offsets $(l_{i,j},t_{i,j},r_{i,j},b_{i,j})$. The final bounding box is obtained by $(x_{i}-l_{i,j},y_{j}-t_{i,j},x_{i}+r_{i,j},y_{j}+b_{i,j})$.
@@ 4.1 Experimental Setup
Model. We present six variants of LoRAT, namely B-224, B-378, L-224, L-378, g-224, and g-378. These variants are based on three sizes of the Vision Transformer (ViT): ViT-B [15], ViT-L [15], and ViT-g [49], serving as the backbone networks. Each size is configured with two distinct input resolutions. Specifically, for the -224 variants, the template size is set to $[112\times 112]$, and the search region size is set to $[224\times 224]$; For the -378 variants, the template size is set to $[196\times 196]$, the search region size is set to $[378\times 378]$. By default, all variants employ the DINO v2[49] pre-trained weights.
Training. The training splits of LaSOT[17], TrackingNet [48], GOT-10k [31] (1k sequences removed for fair comparison following [61, 62], as they overlap with videos used in [33]), and COCO 2017 [40] are used to train the trackers. Additionally, the evaluation results on the GOT-10k test split for trackers are trained only with the GOT-10k training split, following the protocol described in [31]. We train the trackers with 170 epochs, each epoch containing 131,072 image pairs. For the GOT-10k benchmark, we reduce the training epochs to 100 to prevent over-fitting. Each GPU holds 16 samples for an iteration, resulting in a batch size of 128. Following the suggestion in [12], LoRA is applied on all linear layers of ViT backbone, including four projection matrices in the self-attention module and two projection matrices in the MLP module. The rank $r$ of LoRA is set to 64 for all variants. All LoRA linear layers are initialized as the same way as BERT [13] linear layer, truncated normal distribution with std 0.02.
Inference. LoRAT follows conventional steps of Siamese tracking [1] during the inference process: the template is cropped from the first frame of the video sequence, with target object as center; the search area is cropped from the current tracking frame, with the image center as the predicted target center position from the previous frame. To exploit the location prior during tracking, a Hanning window penalty to the classification response map $\mathbf{R}$ output by the head network.
@@
LoRAT-B-224 72.4 81.6 77.9 48.5 61.7 55.3 83.7 88.2 82.2 72.5 82.2 70.6 57.3 58.8
LoRAT-B-378 72.9 82.0 79.1 51.7 65.1 59.2 84.1 88.5 83.2 74.0 83.0 73.2 58.6 60.8
LoRAT-L-224 74.3 83.6 80.8 51.5 65.0 58.7 84.8 89.4 84.3 75.4 84.7 74.9 59.4 61.9
LoRAT-L-378 75.4 84.4 82.2 53.5 67.2 61.4 85.4 89.8 85.6 77.5 86.1 77.6 60.8 63.9
LoRAT-g-224 74.4 83.7 81.4 52.2 60.2 66.1 85.3 90.0 85.1 77.9 88.0 77.9 59.8 62.7
LoRAT-g-378 76.2 85.4 83.5 54.9 69.0 63.2 85.9 90.2 86.3 79.0 87.8 81.0 61.3 64.7
Table 2: Comparison on efficiency with SOTA Transformer trackers. The values in the brackets of # params of our trackers stand for LoRA and extra components (token type embedding and head), respectively.
@@
OSTrack-256 [62] 105.4 21.5 -
OSTrack-384 [62] 58.1 48.3 -
SeqTrack-B256 [6] 40 66 89
SeqTrack-L384 [6] 5 524 309
[…]
ROMTrack [4] 62 34.5 92.1
ROMTrack-384 [4] 28 77.7 92.1

### v2

@@ Abstract
Motivated by the Parameter-Efficient Fine-Tuning (PEFT) in large language models, we propose LoRAT, a method that unveils the power of larger Vision Transformers (ViT) for tracking within laboratory-level resources. The essence of our work lies in adapting LoRA, a technique that fine-tunes a small subset of model parameters without adding inference latency, to the domain of visual tracking. However, unique challenges and potential domain gaps make this transfer not as easy as the first intuition. Firstly, a transformer-based tracker constructs unshared position embedding for template and search image. This poses a challenge for the transfer of LoRA, usually requiring consistency in the design when applied to the pre-trained backbone, to downstream tasks. Secondly, the inductive bias inherent in convolutional heads diminishes the effectiveness of parameter-efficient fine-tuning in tracking models. To overcome these limitations, we first decouple the position embeddings in transformer-based trackers into shared spatial ones and independent type ones. The shared embeddings, which describe the absolute coordinates of multi-resolution images (namely, the template and search images), are inherited from the pre-trained backbones. In contrast, the independent embeddings indicate the sources of each token and are learned from scratch. Furthermore, we design an anchor-free head solely based on a multilayer perceptron (MLP) to adapt PETR, enabling better performance with less computational overhead. With our design, 1) it becomes practical to train trackers with the ViT-g backbone on GPUs with only memory of 25.8GB (batch size of 16); 2) we reduce the training time of the L-224 variant from 35.0 to 10.8 GPU hours; 3) we improve the LaSOT SUC score from 0.703 to 0.742 with the L-224 variant; 4) we fast the inference speed of the L-224 variant from 52 to 119 FPS. Code and models are available at https://github.com/LitingLin/LoRAT.
@@ 1 Introduction
Yet, it is not as easy as our first intuition to apply LoRA to the one-stream tracking architecture, mainly due to two key issues below: ① Ineffective design of positional encoding module. Most Transformer-based trackers adopt separate positional encodings for template and search region tokens. Albeit effective in full fine-tuning [72], such design is not well aligned with LoRA-style PEFT and thus results in suboptimal performance. We argue that this ineffectiveness comes from the disruption to the original structure of the pre-trained model, which is critical to preserve in PEFT, such as LoRA [31]. ② Inductive biases. We empirically identify the convolutional head in the one-stream tracker [72] as a bottleneck in achieving convergence with LoRA. We argue that this is caused by strong inductive biases in convolutional network [16], making it hard to adapt the pre-trained model by fine-tuning a small subset of parameters using LoRA.
Our solution. To overcome the above issues in employing LoRA to one-stream tracker [72], we present two embarrassingly simple but effective designs. Specifically, in addressing issue ①, we draw inspiration from the well-known BERT [14] and introduce the token type embedding in input embedding. Concretely, we assign specific types to tokens of template and search region. By doing so, we can directly leverage existing positional encoding modules of pre-trained ViT models without disrupting their original structure. Considering different resolutions of template and search regions, we further devise a multi-resolution absolute positional embedding adaptation strategy, enabling the support of multi-resolution inputs to pre-trained ViT models for better performance. To deal with issue ②, we propose a multilayer perceptron (MLP)-based anchor-free head network instead of the convolutional head network as in [72], for target classification and bounding box regression. Such a design effectively avoids the inductive biases when applying LoRA to our baseline one-stream tracker, enabling better performance with less computational overhead as shown in our experiments.
It is worth noting that, despite the simplicity in implementation, our solution is specifically devised to employ LoRA better for improving visual tracking. Particularly, with our two designs, we propose LoRAT by applying LoRA to the one-stream OSTrack using various pre-trained ViTs, and extensively assess its performance on five large-scale benchmarks, including LaSOT [18], LaSOText [17], TrackingNet [53], GOT-10k [32], and TNL2K [66]. Our best variant, LoRAT-g-378, sets a new record on LaSOT with a 0.762 SUC score when equipped with ViT-g [54], while the lightest variant, LoRAT-B-224, still achieves a comparable SUC score of 0.717 on LaSOT yet running at 209 FPS. In addition, the required training resources and time are manageable. The training time of different variants, from the light LoRAT-B-224 to strong LoRAT-g-378, ranges from 5.9 to 60 hours using 8 Nvidia V100 GPUs. Notably, our base LoRAT-B-224 can even be trained on a single customer-grade Nvidia RTX 4090 GPU within 11 hours with 605 inference FPS.
@@ 2.1 Paradigms of Transformer Tracker
More recently, the one-stream tracking framework emerges as a significant evolution, with Mixformer [9], SimTrack [6], and OSTrack [72] serving as the vanguard. This paradigm allows for earlier interaction between the template and search region, facilitating joint feature extraction and fusion. The self-attention mechanism of the Transformer architecture plays a crucial role in this context, enhancing the model’s ability to capture complex dependencies. Most subsequent Transformer trackers follow this paradigm. GRM [22] introduces adaptive token division to enable more flexible relation modeling. DropTrack [68] adaptively performs spatial-attention dropout in the frame reconstruction to facilitate temporal correspondence learning in videos. SeqTrack [7] and ARTrack [67] add a Transformer decoder on the top of the one-stream encoder to predict object bounding boxes in an auto-regressive fashion.
@@ 3 Adapting ViT for LoRA-friendly Tracker
Figure 2: Architecture of LoRAT. The template and search region are first split and then projected as patch embeddings. Patch embeddings are added with shared position embeddings and token type embeddings as the input embeddings, which are then fed into Transformer encoder for joint feature extraction and fusion. The resulting representations are fed to the MLP-only head network for target classification and anchor-free-based bounding box regression. Most network components from the pre-trained ViT model are frozen during training, except for LoRA modules applied on the linear layers in the Transformer encoder, the token type embeddings, and the head network.
This section meticulously outlines our approach, beginning with an explanation of our baseline, the one-stream tracking framework. We then introduce Low-Rank Adaptation (LoRA), a technique for efficient fine-tuning of large-scale models for specialized tasks, emphasizing its role in enhancing model adaptability with minimal computational overhead. Following this, we explore the architectural adjustments necessary to make our model compatible with LoRA, including modifications to the model’s input embedding and an MLP-only head network. An overview of our tracker is shown in Figure 2.
@@ 3.1 Preliminaries
One-stream trackers follow the Siamese tracker framework [2], requiring a template image $T\in\mathbb{R}^{H_{T}\times W_{T}\times 3}$ and a search region image $S\in\mathbb{R}^{H_{S}\times W_{S}\times 3}$ as inputs. The template image $T$ and the search region image $S$ are first projected by a learnable linear projection layer to obtain the template token embeddings and the search region token embeddings. Assuming the image patch resolution is $(P,P)$, the template image is divided into $N_{T}=N_{T}^{H}\times N_{T}^{W}$ image patches, and the search region image is divided into $N_{S}=N_{S}^{H}\times N_{S}^{W}$ image patches, where ${N_{T}^{H},N_{T}^{W},N_{S}^{H},N_{S}^{W}}={(H_{T},W_{T},H_{S},W_{S})}/P$. The resulting patch embeddings are the template token embeddings $\mathbf{T}\in\mathbb{R}^{N_{T}\times d}$ and the search region embeddings $\mathbf{S}\in\mathbb{R}^{N_{S}\times d}$, where $d$ represents the hidden dimension of the network.
@@
Low-Rank Adaptation (LoRA) [31] is a technique for efficiently fine-tuning large-scale pre-trained models. Typically, these models are initially trained on extensive datasets, providing broad foundational knowledge. To adapt the models to specific downstream tasks, models are fine-tuned with domain-specific data. Traditional fine-tuning is computationally intensive, as it involves a full forward and backward pass through the network, requiring significant computational resources and memory.
LoRA addresses these issues by proposing a low-rank matrix decomposition to represent weight updates, which can greatly reduce trainable weights and memory requirements. The principle behind LoRA is that while neural networks typically have full-rank weight matrices during pre-training, the updates required for fine-tuning to a specific task have a low “intrinsic dimension”. Remarkably, in many applications, LoRA shows comparable performance to fully fine-tune the network.
@@
Importantly, adopting LoRA does not increase the model’s inference latency. After fine-tuning, the model’s weights are updated by adding the low-rank approximations to the original weights. This approach simplifies the deployment of multiple task-specific models on top of a single large model because the updates ($\Delta\Phi$) are much smaller than the original weights ($\Delta\Theta$).
@@ 3.2 LoRA-friendly Model Design
ViT models require positional encodings to inject positional information [64]. Existing one-stream trackers utilize separate positional embeddings for template and search region tokens, which disrupt the structure of the pre-trained ViT model and lead to ineffectiveness in PEFT. To tackle this, we explore adaptations for multi-resolution positional embeddings and integrate token type embeddings.
@@
Token type embedding. We incorporate token type embedding (also known as segment embedding), originally proposed in the BERT[14] language model, into our tracker assigning unique embeddings for template and search region tokens. This approach decouples token type identification from positional embeddings, improving parameter efficiency and flexibility.
@@
Further, we extend the use of token type embedding to address specific challenges posed by the cropping strategy typically employed in Siamese trackers [2]. The cropping strategy to generate tracking templates has several issues that may confuse the tracker: the aspect ratio of the target objects in the templates is variable; some target objects lack clear semantic meanings; some target objects in the templates can not be clearly distinguished from the background. Token type embedding can help alleviate these issues by explicitly annotating foreground and background parts within the template. We refer to this specific application of token type embedding as foreground object indication embedding.
@@
Positional embedding. Most ViT models employ the 1D absolute positional embedding [16], which only works for fixed resolutions. However, the one-stream trackers usually utilize an image pair of two different sizes as inputs. According to the PEFT-friendly model design principle, sharing the positional embedding across two images is essential. We explore strategies for reusing the standard 1D absolute positional embeddings effectively for different input sizes. Both strategies regard the original 1D absolute embedding in a 2D view. Suppose the original 1D absolute positional embedding is $P$, the 2D view of $P$ can be $P_{2d}=[p_{1,1},...,p_{1,w};p_{2,1},...,p_{2,w};p_{h,1},...,p_{h,w}]$. For convenience, we assume that the resolution of the search region image is the native resolution of the ViT model, thus $P_{2d}$ can be directly applied to search region tokens, i.e. the search region positional embedding $P_{x}=P_{2d}$.
@@
The first approach regards absolute positional embedding as continuous spatial position embedding. We use interpolation to resize $P_{x}$ to the desired resolution for the template, denoted as $P_{z}$. Thus, $P_{z}=\mathrm{interpolate}(P_{x})$. We call this approach as interpolation-based positional embedding adaptation. Notably, this technique is utilized by several Transformer-based trackers, such as OSTrack [72], to initialize distinct positional embeddings for template and search region tokens.
@@
The second approach regards absolute positional embedding as discrete patch index embedding. Thus, with a smaller input resolution, the positional embedding for template tokens $P_{z}$ can be directly taken as a sub-matrix of the $P_{x}$. Specifically, the positional embedding for the template tokens is formed by selecting a block of entries corresponding to the size of the template image patches, $(N_{T}^{H}\times N_{T}^{W})$, from the top left corner of the $P_{x}$. In this way, $P_{z}=[p_{1,1},...,p_{1,N^{W}_{T}};p_{2,1},...,p_{2,N^{W}_{T}};p_{N^{H}_{T},1},...,p_{N^{H}_{T},N^{W}_{T}}]$. We call this approach as slicing-based positional embedding adaptation.
In our ablation experiments, we found the slicing-based adaptation of absolute positional embeddings outperformed the interpolation-based approach. Consequently, we adopted the slicing-based method as our primary strategy for adapting absolute positional embeddings within our tracker.
@@
For other widely utilized positional encoding schemes, such as sinusoidal positional encoding [16], relative positional embedding[59], or Rotary Positional Embedding[60], these do not inherently target specific resolutions and can be directly applied to template tokens and search region tokens.
@@
To mitigate potential inductive biases inherent in convolutional heads, such as locality assumptions on data structure [16], which may impede convergence during LoRA-based fine-tuning, we utilize a multi-layer perceptron (MLP) only head.
The head network generates a redundant set of bounding boxes, each accompanied by a classification score. The classification scores are used to rank the bounding boxes. The bounding box with the highest classification score is considered as the final output. To achieve this, the head network is divided into two branches: one for classification and another for bounding box regression, each consisting of a three-layer MLP. Both branches process the feature map of the search area $\mathcal{S}$, outputted from the Transformer encoder. The head network employs a center-based, anchor-free style [61] to accelerate the training coverage and boost the tracking performance.
@@ 4.1 Experimental Setup
Model. We present six variants of LoRAT, namely B-224, B-378, L-224, L-378, g-224, and g-378. These variants are based on three sizes of the ViT backbone networks: ViT-B [16], ViT-L [16], and ViT-g [54]. Each size is configured with two different input resolutions. For the -224 variants, the template size is set to $[112\times 112]$, and the search region size is set to $[224\times 224]$; For the -378 variants, the template size is set to $[196\times 196]$, the search region size is set to $[378\times 378]$. By default, all variants employ the DINOv2[54] pre-trained weights.
Training. The trackers are trained using the training splits of LaSOT[18], TrackingNet [53], GOT-10k [32] (1k sequences removed for fair comparison following [71, 72], as they overlap with videos used in [35]), and COCO 2017 [43]. For the GOT-10k benchmark, trackers are trained using only the GOT-10k training split, following the protocol in [32]. The trackers are trained for 170 epochs, each epoch containing 131,072 image pairs. For the GOT-10k benchmark, the training epochs are reduced to 100 to mitigate over-fitting. Each GPU holds 16 samples for an iteration, resulting in a batch size of 128. Following the suggestion in [13], LoRA is applied on all linear layers of the ViT backbone, including four projection matrices in the multi-head self-attention module and two projection matrices in the MLP module. The rank $r$ of LoRA is set to 64 for all variants. All LoRA linear layers are initialized using the same method as the BERT [14] linear layer, with a truncated normal distribution with std 0.02.
Inference. LoRAT follows conventional steps of Siamese tracking [2] during inference: the template is cropped from the first frame of the video sequence, centering around the target object; the search area is cropped from the current tracking frame, with the center determined by the bounding box of target object predicted from the previous frame. A Hanning window is added to the classification response map $\mathbf{R}$ to suppress large displacements.
@@
LoRAT-B-224 71.7 80.9 77.3 50.3 61.6 57.1 83.5 87.9 82.1 72.1 81.8 70.7 58.8 61.3
LoRAT-B-378 72.9 81.9 79.1 53.1 64.8 60.6 84.2 88.4 83.0 73.7 82.6 72.9 59.9 63.7
LoRAT-L-224 74.2 83.6 80.9 52.8 64.7 60.0 85.0 89.5 84.4 75.7 84.9 75.0 61.1 65.1
LoRAT-L-378 75.1 84.1 82.0 56.6 69.0 65.1 85.6 89.7 85.4 77.5 86.2 78.1 62.3 67.0
LoRAT-g-224 74.9 84.5 82.3 53.3 65.4 61.1 85.2 89.8 85.1 77.7 87.7 77.7 61.8 66.6
LoRAT-g-378 76.2 85.3 83.5 56.5 69.0 64.9 86.0 90.2 86.1 78.9 87.8 80.7 62.7 67.8
Table 2: Comparison on efficiency with state-of-the-art Transformer trackers. The values in parentheses for the # params of our trackers represent LoRA and extra components (token type embedding and head), respectively. The speed of all trackers was re-evaluated on our machine.
@@
OSTrack-256 [72] 130 21.5 -
OSTrack-384 [72] 68 48.3 -
SeqTrack-B256 [7] 38 66 89
SeqTrack-L384 [7] 6 524 309